prettygirl15:

The ratio of green marbles to yellow marbles is 24 green marbles to 9 yellow marbles. If there are 3 yellow marbles, how many green marbles are there?

prettygirl15:

jhonyy9:

using the rule of 3 simple you can write 24 green to 9 yellow x green to 3 yellow cross multiplie and calcule the x value 9x = 24*3 x = 24*3/9 x = ?